In a proton linear accelerator, protons are accelerated to have a kinetic energy of 600 MeV. What is the speed of these protons? (The rest mass of a proton is 1.67×10−271.67×10−27 kg.) Answer in c unit.
Answer #1

here,

rest mass of proton , m0 = 1.67 * 10^-27 kg

let the speed of these proton be v

the kinetic energy , KE = 600 MeV

KE = 600 * 10^6 * 1.6 * 10^-19 J

m0 * c^2 * ( 1 /sqrt(1 - (v/c)^2) - 1) = 600 * 10^6 * 1.6 * 10^-19 J

1.67 * 10^-27 * (3 * 10^8)^2 * ( 1 /sqrt(1 - (v/(3 * 10^8))^2) - 1) = 600 * 10^6 * 1.6 * 10^-19 J

solving for v

v = 2.38 * 10^8 m/s = 2.38 * 10^8 /( 3 * 10^8) * c = 0.79 c

the speed of these protons is 0.79 c